OpenStudy (anonymous):

lim x-> 4(-) *coming from the left* (√x -2)/(x-4) please show steps

OpenStudy (anonymous):

whats under the sq. rt

OpenStudy (anonymous):

its a the square root of x

myininaya (myininaya):

factor the bottom

myininaya (myininaya):

\[\frac{\sqrt{x}-2}{(\sqrt{x}-2)(\sqrt{x}+2)}\]

myininaya (myininaya):

so we have \[\lim_{x \rightarrow 4^-}\frac{1}{\sqrt{x}+2}=\frac{1}{\sqrt{4}+2}=\frac{1}{4}\]

OpenStudy (anonymous):

oh so you factored the bottom by √x because √x times itself cancels out to be just x right?

myininaya (myininaya):

yes

myininaya (myininaya):

you can also multiply top and bottom by the conjugate of the top

myininaya (myininaya):

this will also work

OpenStudy (anonymous):

can you help me with another one, this one involving piecewise?

myininaya (myininaya):

satellite will help you

OpenStudy (anonymous):

oooh can i write the piecewise function? i mean the latex, not the answer

OpenStudy (anonymous):

lim x->2 where f(x)= { x^2 -4x + 6 X<2 -x^2 +4x-2 x (greater than or equal to) 2

myininaya (myininaya):

plug in 2 into both expressions if you get the same output,then the limit is that if you get different outputs, then the limit is undefined

OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} x^2 + 4x+6 & \text{if} & x <2 \\ - x^2+ 4-2& \text{if} & x \geq 2\ \end{array} \right.\]

OpenStudy (anonymous):

so for the first one x is 18, which doesn't fit with the first set therefore undefined? and the second is 2 which fits so the limit would be 2?

OpenStudy (anonymous):

tada!

OpenStudy (anonymous):

what myininaya said. plug in 2 get different numbers no limit

OpenStudy (anonymous):

oh so the limit would be 2

OpenStudy (zarkon):

yes

myininaya (myininaya):

oh satellite missed an x

myininaya (myininaya):

i will never use satellite's expressions again

myininaya (myininaya):

lol

OpenStudy (zarkon):

and a negative sign

OpenStudy (zarkon):

\[f(x) = \left\{\begin{array}{rcc} x^2 - 4x+6 & \text{if} & x <2 \\ - x^2+ 4x-2& \text{if} & x \geq 2\ \end{array} \right.\]

myininaya (myininaya):

\[2^2-4(2)+6=2=2=-(2)^2+4(2)-2\] so yes the limit is 2

OpenStudy (anonymous):

thank you all. can i ask you all something? how many years of math did you take in college if you did attend?

myininaya (myininaya):

omg several+some more

OpenStudy (zarkon):

10

OpenStudy (zarkon):

4 undergrad...6 grad

OpenStudy (anonymous):

Well I'm trying to advance in math and be as good as you guys

OpenStudy (anonymous):

your help is much appreciated

OpenStudy (anonymous):

i am so ashamed \[f(x) = \left\{\begin{array}{rcc} x^2 -4x+6 & \text{if} & x <2 \\ - x^2+ 4x-2& \text{if} & x \geq 2\ \end{array} \right.\]

OpenStudy (zarkon):

lol

myininaya (myininaya):

oops the units is years sorry 5 to 6 something liek that

OpenStudy (anonymous):

undergrad 0 grad \[\infty\]

myininaya (myininaya):

satellite i was using your function at first

OpenStudy (anonymous):

unfortunately sometimes \[0+\infty=0\] unlike what you may have heard

OpenStudy (anonymous):

yeah i messed up so i take the blame

OpenStudy (zarkon):

must be the 'new math'

OpenStudy (anonymous):

maybe it is \[\frac{0}{\infty}\]

myininaya (myininaya):

i can't read this before tomorrow and know what i'm talking about i think i will go to sleep

OpenStudy (zarkon):

night

myininaya (myininaya):

i'm reading something i know nothing about and my adviser wants me lead a discussion on it tomorrow

OpenStudy (zarkon):

good luck!

myininaya (myininaya):

at least its informal

OpenStudy (zarkon):

that's good

myininaya (myininaya):

goodnight guys

OpenStudy (anonymous):

one more question im doing this continuity and discontinuity at points and which are the points where it isnt continuous. it says which of the discontinuities are removable? what exactly does that mean by the question?